# Avoiding the Price of Adaptivity: Inference in Linear Contextual Bandits via Stability

### v1

@@ 7 Conclusion
We studied statistical inference in linear contextual bandits under adaptive data collection, with a focus on the validity of classical Wald-type confidence intervals. By introducing a penalized variant of the EXP4 algorithm, we demonstrated that it is possible to simultaneously enforce the Lai–Wei stability condition and achieve regret guarantees that are minimax optimal up to logarithmic factors. As a consequence, ordinary least-squares estimators satisfy a central limit theorem, and Wald confidence intervals for linear functionals are asymptotically valid without incurring the $\sqrt{d\log T}$ price of adaptivity. Together, these results establish that stability and statistical efficiency are not fundamentally incompatible in contextual bandit problems, and that careful algorithmic design can recover classical inferential guarantees even under adaptive sampling.
@@ Appendix A Proof of important results
A.1 Proof of Lemma 3
We use the shorthand $\widetilde{g}_{t}=\widehat{g}_{t}+\lambda\nabla R(w_{t})$ as the gradient estimate of the penalized loss function at $w_{t}$. We have
@@ A.1 Proof of Lemma 3
Furthermore, the coordinate-wise positivity of $\widetilde{g}_{t}$ in (48) yields $\nabla\phi(w^{+}_{t+1})\leq\nabla\phi(w_{t})$. Hence, by applying the second order Taylor expansion in equation (A.1), we have that for an intermediate point $z_{t}=\delta\ \nabla\phi(w^{\star}_{t+1})+(1-\delta)\nabla\phi(w_{t})$,

### v2

@@ 7 Conclusion
We studied statistical inference in linear contextual bandits under adaptive data collection, with a focus on the validity of classical Wald-type confidence intervals. By introducing a regularized variant of the EXP4 algorithm, we demonstrated that it is possible to simultaneously enforce the Lai–Wei stability condition and achieve regret guarantees that are minimax optimal up to logarithmic factors. As a consequence, ordinary least-squares estimators satisfy a central limit theorem, and Wald confidence intervals for linear functionals are asymptotically valid without incurring the $\sqrt{d\log T}$ price of adaptivity. Beyond asymptotic validity, our analysis provides explicit rates of convergence to normality, offering a quantitative characterization of how stability, regret, and inferential accuracy interact in finite samples. As a consequence of our results we obtain valid Wald-type inference for conditional average treatment effects in linear models, even when treatments are assigned adaptively via contextual bandit algorithms. Together, these results establish that stability and statistical efficiency are not fundamentally incompatible in contextual bandit problems, and that careful algorithmic design can recover classical inferential guarantees even under adaptive sampling.
@@ Appendix A Proof of important results
Proof of Lemma 6
We use the shorthand $\widetilde{g}_{t}=\widehat{g}_{t}+\lambda\nabla R(w_{t})$ as the gradient estimate of the regularized loss function at $w_{t}$. We have
@@
Furthermore, as $\ell_{t}\geq 0$ (Assumption (A3)) the coordinate-wise positivity of $\widetilde{g}_{t}$ in (61) yields $\nabla\phi(w^{+}_{t+1})\leq\nabla\phi(w_{t})$. Hence, by applying the second order Taylor expansion in equation (A), we have that for an intermediate point $z_{t}=\delta\ \nabla\phi(w^{\star}_{t+1})+(1-\delta)\nabla\phi(w_{t})$,